OpenStudy (anonymous):

how do you solve x+4/2=7

OpenStudy (anonymous):

multiply by 2 gives u x+4=14 then subtract 4 to each side gives u 10 so x=10 thats how i do it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

so to solve x+20=3 would be the same /2 The x and the 20 upunder 2 is the same

OpenStudy (anonymous):

ow wait the equation is X/2+20=3

OpenStudy (anonymous):

x+20 all divied by two =3

OpenStudy (anonymous):

ow k so u multiply by 2 gives you x+20 = 6 then -20 -20 ------------ the 20's cancel then 6 -20 = -14 so x equals -14

OpenStudy (anonymous):

say ur equation is useing subtraction you add to get watever number away from x. if its multiplication u divide and the other way around for both so example > 4x-20=80 divide by 4 to both sides then the 4s cancel so x-20 =20 its subtraction now so you +20 to both sides 20's on the left cancel then you got x=40